# Interlayer interactions in $\text{La}_{3}\text{Ni}_{2}\text{O}_{7}$ under pressure: from $s^{\pm}$ to $d_{xy}$-wave superconductivity

### v1

@@ Abstract
Superconducting symmetries.— For a given set of parameters entering $\bar{\Gamma}(\bm{k},\bm{k}^{\prime})$, the solution of Eq. (7) with the highest eigenvalue $\lambda_{\rm max}$ will correspond to the highest critical temperature $T_{c}$ and the symmetry of the pairing gap is reflected in the corresponding eigenfunction $\Delta_{\rm max}(\bm{k})$. In the following, we consider the reference value of $U\approx 1.16$ eV, which has been estimated to match $T_{c}\approx 80$ K [12]. Considering the largest interlayer hopping in the model ($|t_{\perp}^{z}|\approx 0.635$ eV) [6], this gives a moderate $U/|t_{\perp}^{z}|\approx 1.83$, reflecting a itinerant electron picture. For these parameters $U_{\perp}\!=\!0$ the leading and subleading solutions of Eq. (7) display $s^{\pm}$-wave and $d_{xy}$-wave symmetries (with $\lambda^{s^{\pm}}/\lambda^{d_{xy}}>1$), in accordance with previous studies [11, 14, 8].
@@
For increasing values of $U_{\perp}$, however, the ratio $\lambda^{s^{\pm}}/\lambda^{d_{xy}}$ decreases and eventually the solution with $d_{xy}$-wave symmetry becomes dominant. This is illustrated in Figure 2(a), which shows a color map of $\log{(\lambda^{s^{\pm}}/\lambda^{d_{xy}})}$ for different values of onsite $U$ and interlayer $U_{\perp}$ interaction strengths. The dashed line represents the curve $\lambda^{s^{\pm}}=\lambda^{d_{xy}}$ and can be interpreted as a “phase boundary” separating the regions where each gap symmetry dominates.
@@
For the reference value of $U\!=\!1.16$ eV, we find that the transition point occurs at a moderate value of $U^{c}_{\perp}\approx 0.2U=0.23$ eV, with the gap functions displaying $s^{\pm}$-wave and $d_{xy}$-wave symmetries for $U_{\perp}<U^{c}_{\perp}$ and $U_{\perp}>U^{c}_{\perp}$ (Figs. 2(b) and (c)), respectively. For larger values of $U$, however, the $s^{\pm}$-wave symmetry is clearly favored even for ratios $U_{\perp}/U$ as large as $50-60\%$.
@@
Some insight as to why the interlayer interaction tends to favor a d-wave symmetry over the $s^{\pm}$ one can be gained by looking more closely on the role of the interorbital contribution to the pairing interaction. Imagine if one could tune the relative contributions of the different FS sheets to the paring gap function defined by Eq. (7). In practice, this can be done by artificially giving different weights to the terms in the sum on the left-hand-side of Eq. (7). For instance, one could exclude (assign zero weight to) the contribution of one of the FS pockets ($\alpha$, $\beta$, or $\gamma$) and see what the resulting symmetry of the leading pairing gap function looks like.
This is done in Fig. 3 for $U=1.16$ eV. Figure 3(a) shows that, for $U_{\perp}=0$, the paring gap symmetry remains $s^{\pm}$-wave even if the contribution of the $\gamma$ pocket is excluded and only pockets $\alpha$ and $\beta$, which are dominated by $3d_{x^{2}-y^{2}}$ orbitals and hence have a more “planar” character, are considered in the calculation. In a sense, this would mimic a “cuprate-like” scenario [16]. In this situation, by increasing $U_{\perp}$ the symmetry of the leading gap function becomes $d_{x^{2}-y^{2}}$ and not $d_{xy}$, as in the case where all pockets are considered (Fig. 2(c)). One the other hand, the pairing symmetry becomes $d_{xy}$ if, instead, the contribution of the $\alpha$ pocket is excluded (Fig. 3(c)). Even though Fig. 3(c) shows results for $U_{\perp}=0$, we have checked that this result holds for a large range of $U_{\perp}$ values. Finally, if the contribution of the $\beta$ pocket is excluded, the gap symmetry becomes $d_{x^{2}-y^{2}}$ (Fig. 3(d)) and not $d_{xy}$.
These results indicate that, when all three pockets are considered in the calculation (Fig. 2), by increasing $U_{\perp}$, the relative contribution of the $\gamma$ pocket (dominated by $3d_{z^{2}}$ orbitals) increases over the (more planar) $\alpha$ pocket contribution, thereby inducing the switching between $s^{\pm}$- and nodal $d_{xy}$-wave paring symmetries. As such, the interplay between $\gamma$ and $\alpha$ pockets in the pairing vertex is a crucial element in explaining the $s^{\pm}\rightarrow d_{xy}$ order switching as $U_{\perp}$ increases.
Figure 4: Enhancement of the pairing vertex function ($\Gamma(\bm{k},\bm{k}^{\prime})$) by interlayer interactions. Panels (a) and (b) show the ratio of $\Gamma(\bm{k},\bm{k}^{\prime})$ calculated with $U_{\perp}=0.348$ eV and $\Gamma^{\mathrm{on}}(\bm{k},\bm{k}^{\prime})$, calculated with $U_{\perp}=0$ in a logarithmic scale for a fixed $\bm{k}$ (black ring) located at (a) $\gamma$ and (b) $\alpha$ Fermi surface pockets, dominated by $d_{3z^{2}-r^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals respectively.
@@
For example, as shown in Fig., 4(a), for a $\bm{k}$-point located in the $\gamma$ pocket, increasing $U_{\perp}$ leads to an enhancement of the pairing vertex function ($\Gamma(\bm{k},\bm{k}^{\prime})>\Gamma^{\rm on}(\bm{k},\bm{k}^{\prime})$) along the $\alpha$ pocket and a decrease ($\Gamma(\bm{k},\bm{k}^{\prime})<\Gamma^{\rm on}(\bm{k},\bm{k}^{\prime})$) along the $\beta$ pocket. A similar result occurs when $\bm{k}$ is located at the $\alpha$ pocket (Fig, 4(b)): the pairing vertex function is enhanced along the $\gamma$ pocket and decreases along the $\beta$ pocket.
@@
Concluding remarks.— In this work, we investigate the role of interlayer Coulomb interaction terms on the superconducting properties of $\text{La}_{3}\text{Ni}_{2}\text{O}_{7}$ under pressure. Our results indicate that for moderate onsite interactions in the model (up to $U\sim 1.1$ eV), and relatively small interlayer interaction strengths (about $\sim 30\%$ of the onsite term) the leading pairing symmetry becomes of the $d_{xy}$-wave type instead of the $s^{\pm}$-wave predicted by several works in the absence of interlayer interactions [11, 14, 16, 15, 11, 12, 31, 32, 8]. Our calculations also confirm that the $s^{\pm}$ solution is dominant for large onsite interactions ($U\gtrsim 1.3$ eV), which agrees with previous calculations based on $t-J-J_{\perp}$ models [14].
@@
More interestingly, our analysis shows that the superconducting pairing symmetry is strongly dependent on the interplay of all three Fermi surface pockets. In particular, the interlayer interaction tends to favor interorbital pairing between electrons with $d_{3z^{2}-r^{2}}$ and $d_{x^{2}-y^{2}}$ orbital character, underscoring the view that a complete description of the superconducting properties can only be achieved by multi-orbital models [8, 10, 16, 9].
@@
We acknowledge financial support from the São Paulo Research Foundation (FAPESP), Brazil (process numbers 2022/15453-0 and 2023/14902-8) and from CNPq (Grant No. 312622/2023-6).
S1 Matrix random-phase approximation
Since our investigation focuses on exploring the role of multiorbital effects on the magnetic fluctuations of the pressurized bilayer nickelate, we consider the two-orbital $d_{x^{2}-y^{2}}$ and $d_{z^{2}-r^{2}}$ low-energy model [6], as represented in Fig. 1(a) of the main text. The non-interacting Hamiltonian reads
$\begin{split}H_{0}&=\sum_{\bm{k}\sigma sS}(T_{\bm{k}}^{s}-\mu)c_{\bm{k}sS\sigma}^{\dagger}c_{\bm{k}sS\sigma}\\ &+\sum_{\bm{k}\sigma s}t_{\perp}^{s}(c_{\bm{k}sA\sigma}^{\dagger}c_{\bm{k}sB\sigma}+\text{h.c.})\\ &+\sum_{\bm{k}\sigma(s\neq p)S}V_{\bm{k}}c_{\bm{k}sS\sigma}^{\dagger}c_{\bm{k}pS\sigma}\\ &+\sum_{\bm{k}\sigma(s\neq p)}V_{\bm{k}}^{\prime}(c_{\bm{k}sA\sigma}^{\dagger}c_{\bm{k}pB\sigma}+\text{h.c.}),\end{split}$ (S1)
where $c^{\dagger}_{\bm{k}sS\sigma}$ creates an electron with spin $\sigma$, in the orbital $s=\{x,z\}$ corresponding to $\{d_{x^{2}-y^{2}},d_{3z^{2}-r^{2}}\}$, sublattice $S=\{A,B\}$, and momentum $\bm{k}$, $\mu$ is the chemical potential, and h.c. shortcuts Hermitian conjugate. This model considers up to second-neighbor hoppings such that
$\begin{split}T_{\bm{k}}^{s}&=2t_{1}^{s}(\cos k_{x}+\cos k_{y})+4t_{2}^{s}\cos k_{x}\cos k_{y}+\epsilon^{s}\\ V_{\bm{k}}&=2t_{3}^{xz}(\cos k_{x}-\cos k_{y}),\\ V_{\bm{k}}^{\prime}&=2t_{4}^{xz}(\cos k_{x}-\cos k_{y}).\end{split}$ (S2)
The hopping parameters are given in Ref. [6]. Fig. 1(c) of the main text shows the band structure with its respective orbital composition. There is an interplay between the two orbitals $d_{x^{2}-y^{2}}$ (red) and $d_{3z^{2}-r^{2}}$ (blue) along the path $\bm{\Gamma}-\bm{X}-\bm{M}$. Also, the $\gamma$ pocket is dominated by the $d_{3z^{2}-r^{2}}$ orbital and contributed by a nearly flat hole pocket centered at the $\bm{M}$ point.
We consider interactions through a degenerate Hubbard model including interlayer interactions. Within this model, the electrons are not constrained to be localized and the specific values of the interaction strengths are not restricted to the limit $U/t\gg 1$ such as the $t-J$-like models considered in the previous literature. In fact, it has been estimated using the random-phase approximation within mean-field theory that $U\approx 1.16$ eV to match $T_{c}=80$ K [liu\$s^ifmmodepmelsetextpmfi\$WavePairingDestructive2023], while the largest hopping is $t_{\perp}^{z}\approx-0.635$ eV [6] such that $U/t_{\perp}^{z}\approx-1.83$, which is a moderate value. The interaction Hamiltonian for the two-orbital itinerant electron model is given in the main text. Since we are interested in the magnetic fluctuations caused by interacting electrons, we focus our study on the bare and interacting susceptibilities. Having in mind that the present non-interacting model [Eq. (S1)] is independent of the spin directions, the bare susceptibility is the same for charge or spin channels and can be defined as [19]
[…]
where $i\nu_{m}$ denotes bosonic frequencies, $\tau$ imaginary time, $T$ temperature, and $T_{\tau}$ time-ordering. We compute the bare susceptibility in the Green’s function form performing the Matsubara summation numerically using Ozaki’s method [33]. The formalism reads
$\begin{split}&[\hat{\chi}]^{pAqB}_{sCtD}(\bm{q},i\nu_{m})=-\frac{T}{N}\sum_{\bm{k}i\omega_{n}}G_{sCpA}(\bm{k},i\omega_{n})G_{qBtD}(\bm{k}+\bm{q},i\omega_{n}+i\nu_{m}),\end{split}$ (S4)
where
$G_{sCpA}(\bm{k},i\omega_{n})=\sum_{\nu_{\bm{k}}}\frac{\psi^{sC}_{\nu_{\bm{k}}}(\bm{k})\psi^{pA*}_{\nu_{\bm{k}}}(\bm{k})}{i\omega_{n}-E_{\nu_{\bm{k}}}(\bm{k})},$ (S5)
and $\psi^{sC}_{\nu_{\bm{k}}}(\bm{k})$ are non-interacting electron eigenfunctions related to the $\nu_{\bm{k}}$-th $E_{\nu_{\bm{k}}}(\bm{k})$ energy band of Eq. (S1). We adopt $T=0.02$ eV.
$\hat{\chi}_{s}(\bm{q},i\nu_{m})$ $=\hat{\chi}(\bm{q},i\nu_{m})[\hat{1}-\hat{U}_{s}\hat{\chi}(\bm{q},i\nu_{m})]^{-1},$ (S6)
$\hat{\chi}_{c}(\bm{q},i\nu_{m})$ $=\hat{\chi}(\bm{q},i\nu_{m})[\hat{1}+\hat{U}_{c}\hat{\chi}(\bm{q},i\nu_{m})]^{-1}.$ (S7)
S2 Pairing symmetry of individual pockets

### v3

@@ IV.1 Leading solutions of the gap equation
In the following, we consider the reference value of $U\approx 1.16$ eV, which has been estimated as $T_{c}\approx 80$ K from real-space mean-field calculations [12]. Considering the largest interlayer hopping in the model ($|t_{\perp}^{z}|\approx 0.635$ eV) [6], gives a moderate $U/|t_{\perp}^{z}|\approx 1.83$, reflecting an itinerant electron picture. For these parameters and $U_{\perp}\!=\!0$, the leading and subleading solutions of Eq. (14) display $s^{\pm}$-wave and $d_{xy}$-wave symmetries (with $\lambda^{s^{\pm}}/\lambda^{d_{xy}}>1$), in accordance with previous studies [11, 14, 8].
@@
For increasing values of $U_{\perp}$, however, the ratio $\lambda^{s^{\pm}}/\lambda^{d_{xy}}$ decreases and eventually the solution with $d_{xy}$-wave symmetry becomes dominant. This is illustrated in Figure 3(a), which shows a color map of $\log{(\lambda^{s^{\pm}}/\lambda^{d_{xy}})}$ for different values of onsite $U$ and interlayer $U_{\perp}$ interaction strengths. The dashed line represents the curve $\lambda^{s^{\pm}}=\lambda^{d_{xy}}$ and can be interpreted as a “phase boundary” separating the regions where each gap symmetry dominates. The corresponding superconducting gap symmetry solutions projected onto the Fermi surface for the dominant $s^{\pm}$ and $d_{x^{2}-y^{2}}$ waves are shown in Fig. 3(b) and (c), respectively.
Figure 4: Line-cuts of the phase diagram shown in Fig. 3 showing the leading and subleading eigenvalues of Eq. (14) as a function of (a,b,c) $U$ for different values of $U_{\perp}/U$ and (d) $U_{\perp}/U$ for $U=1.16$ eV.
A more detailed view of the leading and sub-leading solutions as a function of $U$ and $U_{\perp}$ is shown in Fig. 4. For $U/U_{\perp}=0$ [Fig. 4(a)], $s^{\pm}$ and $d_{x^{2}-y^{2}}$ waves are nearly degenerate for $U<1$ eV, while the $s^{\pm}$ dominates for larger values of $U$. By contrast, increasing the ratio $U_{\perp}/U>0$ immediately breaks the degeneracy and suppresses the $s^{\pm}$ solution, although higher $U$ again favor the sign-changing $s$-wave, as shown in Figs. 4(b) and (c).
@@
For the reference value of $U\!=\!1.16$ eV, we show in Fig. 4(d) that the transition point occurs at a moderate value of $U^{c}_{\perp}\approx 0.2U=0.23$ eV, with the gap functions displaying $s^{\pm}$-wave and $d_{xy}$-wave symmetries for $U_{\perp}<U^{c}_{\perp}$ and $U_{\perp}>U^{c}_{\perp}$ [Figs. 3(b) and (c)], respectively. For larger values of $U$, however, the $s^{\pm}$-wave symmetry is clearly favored even for ratios $U_{\perp}/U$ as large as $50-60\%$.
Additionally, as discussed in Appendix A, we note that the Hund’s couplings $J$ and $J_{\perp}$ only marginally affect this general behavior. In fact, increasing $J/U$ does not influence the leading pairing symmetry [Fig. 8]. Moreover, we find that changing $J_{\perp}/U_{\perp}$ does not change our conclusions qualitatively, although decreasing $J_{\perp}/U_{\perp}$ shifts up the phase boundary separating the two symmetries [dashed black line in Fig. 3(a)].
IV.2 The role of the charge susceptibility
A key aspect in the transition of $s^{\pm}$ to $d_{xy}$ pairing symmetries as a function of $U_{\perp}$ is its connection the to the decrease in the charge susceptibility peaks near the nesting vectors, previously shown in Fig. 2. As previously mentioned, Eq. (6) shows that the spin susceptibility is unaffected by the interlayer interactions. On the other hand, Figs. 2(c,d) show that the charge susceptibility is suppressed for increasing $U_{\perp}/U$ in regions around the $\bm{\Gamma}$ and $\bm{M}$ pockets.
This is better illustrated in Figs. 5(a) and (b), which show the main eigenvalue of the spin [$\hat{\chi}_{s}(\bm{q},0)$] and charge [$\hat{\chi}_{c}(\bm{q},0)$] susceptibility matrices respectively along the $\bm{\Gamma}-\bm{X}-\bm{M}-\bm{\Gamma}$ high-symmetry path. The main three nesting vectors ($\bm{q}_{1}$, $\bm{q}_{2}$ and $\bm{q}_{3}$) are marked by small colored arrows.
Figures 5(c,d) show the same gap function data as Fig. 1(b,c) [$U_{\perp}/U=0.1$ ($\triangle$) and $U_{\perp}/U=0.3$ ($\square$), respectively] with a shift in the $k_{x}$ component so that the three nesting vectors $\bm{q}_{i=1,2,3}$ (represented as colored arrows, in scale) can be better visualized. A comparison between Figs. 5(c) and (d) shows that the $\bm{q}_{1}$ and $\bm{q}_{2}$ nesting vectors connect pockets where the leading gap function has the same sign in the $s^{\pm}$-wave case (i.e., $\Delta^{\alpha}(\bm{k})\!=\!+\Delta^{\alpha}(\bm{k}+\bm{q}_{i=1,2})$) while they are of opposite sign in the $d_{xy}$-wave case ($\Delta^{\alpha}(\bm{k})\!=\!-\Delta^{\alpha}(\bm{k}+\bm{q}_{i=1,2})$).
These results can be better understood by looking at both the dependence of the pairing vertex function on the charge susceptibility [Eq. (12)] and on the pairing function $\bar{\Gamma}(\bm{k},\bm{k}^{\prime})$ on the left-hand-side of the linear pairing equation [Eq. (14)]. First, we remark that $\hat{\chi}_{s}(\bm{q},0)$ contributes with a positive sign to the right-hand-side of Eq. (12), while $\hat{\chi}_{c}(\bm{q},0)$ contributes with a negative sign. As such, reductions in $\hat{\chi}_{c}(\bm{q},0)$ tend to increase the pairing $\bar{\Gamma}(\bm{k},\bm{k}+\bm{q})$, particularly at the nesting points $\bm{q}=\bm{q}_{1,2}$, at which the spin susceptibility is lower than at $\bm{q}=\bm{q}_{3}$ (see Fig. 5(a)).
Due to negative sign in the left-hand side in Eq. (14) and considering that $\bar{\Gamma}(\bm{k},\bm{k}+\bm{q})$ is positive at the nesting points, we argue that such increases in $\bar{\Gamma}(\bm{k},\bm{k}+\bm{q}_{1,2})$ as $U_{\perp}/U$ increases will favor sign-changing solutions which obey $\Delta^{\alpha}(\bm{k})\!=\!-\Delta^{\alpha}(\bm{k}+\bm{q}_{1,2})$ such as the $d_{xy}$-wave symmetric one.
Figure 5: Susceptibilities and the sign changes in the superconducting gap. The main eigenvalue of the spin and charge susceptibility matrices are shown in panels (a) and (b), respectively. Dashed lines represent results for different $U_{\perp}/U$ ratios, while the arrows point out to the susceptibility peaks related to each putative nesting vector $\bm{q}_{i}$. Panels (c) and (d) show the same gap function data of Fig. 1(b) and (c), respectively, while the colored arrows corresponded to the vectors $\bm{q}_{i}$ in the upper panels.
IV.3 Orbital contributions to the pairing interaction
Further insight as to why the interlayer interaction tends to favor a $d$-wave symmetry over the $s^{\pm}$ one can be gained by looking more closely on the role of the interorbital contribution to the pairing interaction. Imagine if one could tune the relative contributions of the different FS sheets to the paring gap function defined by Eq. (14). In practice, this can be accomplished by artificially giving different weights to the terms in the sum on the left-hand-side of Eq. (14). For instance, one could exclude (assign zero weight to) the contribution of one of the FS pockets ($\alpha$, $\beta$, or $\gamma$) and see what the resulting symmetry of the leading pairing gap function looks like.
@@ IV.3 Orbital contributions to the pairing interaction
Table 1: Summary of superconducting pairing symmetries found for each pocket or combination of pockets $\alpha,\beta,\gamma$ as indicated in Fig. 1(b). We set $U=1.16$ eV.
Pockets $U_{\perp}/U$ Symmetry
$\alpha$ any $g$-wave
$\beta$ any $d_{xy}$-wave
$\gamma$ any $d_{x^{2}-y^{2}}$-wave
$\alpha+\beta$ $\lesssim 0.3$ $s^{\pm}$-wave
$\alpha+\beta$ $\gtrsim 0.3$ $d_{xy}$-wave
$\alpha+\gamma$ any $d_{x^{2}-y^{2}}$-wave
$\beta+\gamma$ any $d_{xy}$-wave
all $\lesssim 0.2$ $s^{\pm}$-wave
all $\gtrsim 0.2$ $d_{xy}$-wave
This is done in Fig. 6 for $U=1.16$ eV. Figure 6(a) shows that, for $U_{\perp}=0$, the paring gap symmetry remains $s^{\pm}$-wave even if the contribution of the $\gamma$ pocket is excluded and only pockets $\alpha$ and $\beta$, which are dominated by $3d_{x^{2}-y^{2}}$ orbitals and hence have a more “planar” character, are considered in the calculation. In a sense, this would mimic a “cuprate-like” scenario [17]. In this situation, by increasing $U_{\perp}$ the symmetry of the leading gap function becomes $d_{x^{2}-y^{2}}$ and not $d_{xy}$, as in the case where all pockets are considered [Fig. 3(c)]. On the other hand, the pairing symmetry becomes $d_{xy}$ if, instead, the contribution of the $\alpha$ pocket is excluded [Fig. 6(c)]. Finally, if the contribution of the $\beta$ pocket is excluded, the gap symmetry becomes $d_{x^{2}-y^{2}}$ [Fig. 6(d)] and not $d_{xy}$. A summary of these findings is presented in Table 1, which also includes other combinations of the $\alpha,\beta,\gamma$ pockets (as detailed in Appendix B).
These results indicate that, when all three pockets are considered in the calculation (Fig. 3), by increasing $U_{\perp}$, the relative contribution of the $\gamma$ pocket (dominated by $d_{3z^{2}-r^{2}}$ orbitals) increases over the (more planar) $\alpha$ pocket contribution, thereby inducing the switching between $s^{\pm}$- and nodal $d_{xy}$-wave paring symmetries. As such, the interplay between $\gamma$ and $\alpha$ pockets in the pairing vertex is a crucial element in explaining the $s^{\pm}\rightarrow d_{xy}$ order switching as $U_{\perp}$ increases.
@@
Figure 7: Enhancement of the pairing vertex function ($\Gamma(\bm{k},\bm{k}^{\prime})$) by interlayer interactions. Panels (a) and (b) show the ratio of $\Gamma(\bm{k},\bm{k}^{\prime})$ calculated with $U_{\perp}=0.348$ eV and $\Gamma^{\mathrm{on}}(\bm{k},\bm{k}^{\prime})$, calculated with $U_{\perp}=0$ in a logarithmic scale for a fixed $\bm{k}$ (black ring) located at (a) $\gamma$ and (b) $\alpha$ FS pockets, dominated by $d_{3z^{2}-r^{2}}$ and $d_{x^{2}-y^{2}}$ orbitals respectively.
For example, as shown in Fig. 7(a), for a $\bm{k}$-point located in the $\gamma$ pocket, increasing $U_{\perp}$ leads to an enhancement of the pairing vertex function [$\Gamma(\bm{k},\bm{k}^{\prime})>\Gamma^{\rm on}(\bm{k},\bm{k}^{\prime})$] along the $\alpha$ pocket and a decrease [$\Gamma(\bm{k},\bm{k}^{\prime})<\Gamma^{\rm on}(\bm{k},\bm{k}^{\prime})$] along the $\beta$ pocket. A similar result occurs when $\bm{k}$ is located at the $\alpha$ pocket [Fig, 7(b)]: the pairing vertex function is enhanced along the $\gamma$ pocket and decreases along the $\beta$ pocket.
@@ V Concluding remarks
In this work, we investigate the role of interlayer Coulomb interaction terms on the superconducting properties of $\text{La}_{3}\text{Ni}_{2}\text{O}_{7}$ under pressure in the weak- to intermediate-coupling limit. Our results indicate that for moderate onsite interactions in the model (up to $U\sim 1.1$ eV), and relatively small interlayer interaction strengths (about $\sim 30\%$ of the onsite term) the leading pairing symmetry becomes of the $d_{xy}$-wave type instead of the $s^{\pm}$-wave predicted by several works in the absence of interlayer interactions [8, 12, 8, 14, 15]. Our calculations also confirm that the $s^{\pm}$ solution is dominant for large onsite interactions ($U\gtrsim 1.3$ eV). This is consistent with results from strong-coupling ($t\!-\!\tilde{J}\!-\!\tilde{J}_{\perp}$) bi-layer single-orbital models which show $s^{\pm}$-wave superconductivity favored by a strong interlayer exchange coupling [14, 15].
Additionally, our results indicate that charge fluctuations play an essential role in the $s$- to $-d$-wave transition as the interlayer interaction $U_{\perp}$ increases. This is a consequence of a decrease in the charge susceptibility for large $U_{\perp}$ values, particularly at momenta $\bm{q}_{i}$ satisfying the nesting condition between different FS sheets. This in turn leads to an increase of the RPA paring function at these nesting vectors, favoring $d_{xy}$-wave sign-changing solutions obeying $\Delta^{\alpha}(\bm{k})\!=\!-\Delta^{\alpha}(\bm{k}+\bm{q}_{i})$ over the $s^{\pm}$-wave ones. We note that these $s^{\pm}$-wave solutions differ from those appearing in iron-based superconductors [23, 24], for which sign changes can be traced back to the spin susceptibility at the nesting points.
Figure 8: Leading gap functions $\Delta_{\rm max}(\bm{k})$ displaying $s^{\pm}$-wave and $d_{xy}$-wave symmetries for $U=1.16$ eV and $U_{\perp}/U=0.1$ ($\triangle$) and $U_{\perp}/U=0.3$ ($\square$).
@@
More interestingly, our analysis shows that the superconducting pairing symmetry is strongly dependent on the interplay of all three FS pockets. In particular, the interlayer interaction tends to favor interorbital pairing between electrons with $d_{3z^{2}-r^{2}}$ and $d_{x^{2}-y^{2}}$ orbital character, underscoring the view that a complete description of the superconducting properties can only be achieved by multi-orbital models [8, 10, 17, 9]. In fact, we expect that a similar picture can emerge in systems exhibiting multiple FS pockets with close competing pairing symmetries such as the three-layer La4Ni3O10 [32].
Note added: After the submission of this manuscript, we became aware of related works [35, 36] reporting transitions from $s$-wave to $d$-wave pairing symmetry as a function of the interlayer interaction strength. Although these works use different methods (namely FLEX [35] and fRG [36]), the main results are qualitatively similar to our findings.
VI Acknowledgments
@@
We thank Ilya Eremin and Steffen Bötzel for fruitful discussions. We acknowledge financial support from the São Paulo Research Foundation (FAPESP), Brazil (2022/15453-0 and 2023/14902-8) and from CNPq (Grant No. 312622/2023-6).
VII Data availability
The data that support the findings of this paper are openly available on Zenodo [37].
Figure 9: Same as Fig. 6 but for individual pockets $\beta$ (a,b), $\gamma$ (c), and $\alpha$ (d).
Appendix A Influence of Hund’s couplings in the pairing symmetry
The influence of Hund’s couplings $J$ and $J_{\perp}$ in the pairing symmetry is shown in Fig. 8. For $U/U_{\perp}=0$, Fig. 8(a) shows that $J=J_{\perp}=0$ results in a leading $s^{\pm}$-wave, while increasing $J/U$ does not influence the pairing symmetry but for enhancing the gap function at the $\alpha$ pocket in comparison with the other pockets, which is illustrated in panel (b). On the other hand, fixing $J=U/4$ and $J_{\perp}=0$, the $U_{\perp}/U=0.3$ gap symmetry shown in Fig. 8(c) results in a leading $s^{\pm}$-wave, while only further increase of interlayer interactions, $U_{\perp}/U=0.5$ as in Fig. 8(d), induces the subleading $d_{xy}$ as the leading pairing symmetry. Therefore, lowering $J_{\perp}/U_{\perp}$ only increases the ratio $U_{\perp}/U$ necessary for the $s^{\pm}\rightarrow d_{xy}$ transition, which does not qualitatively change the phase diagram of Fig. 3(a).
Appendix B Pairing symmetry of individual pockets